On the Topography of Maass Waveforms for PSL(2, Z)

Dennis A. Hejhal Barry N. Rackner

Abstract

This article provides a glimpse into “arithmetical quantum chaos” through a study of the topography and statistical properties of the eigenfunctions of the Laplacian for the modular surface mathrmP SL(2, ℤ)\H.